# Adaptive Cohen’s Class Time-Frequency Distribution

### v1

@@ IV Numerical experiments
Red noise Margenau-Hill $2.3888$ $6.3998$ $3.6360$ $5.2562$
Kirkwood-Rihaczek $1.7774$ $5.9456$ $3.4881$ $5.1559$
Born-Jordan $4.0111$ $7.4167$ $4.3805$ $6.6348$
Page $1.3606$ $5.9863$ $3.0925$ $3.8628$
Wiener filter $11.8832$ $\mathbf{16.7052}$ $\mathbf{13.1156}$ $\mathbf{11.9071}$
Adaptive CCTFD $\mathbf{20.4654}$ $16.1495$ $7.0496$ $7.2056$
V Conclusion
The convolution type of CCTFD time-frequency analysis method suitable for the adaptive filter denoising for additive noises jamming signals under the condition of low SNR was established, giving rise to the adaptive kernel function which minimizes the MSE in the WVD domain. The proposed adaptive CCTFD can automatically adjust its kernel function according to the change of signal to adapt to different signal characteristics. It turns out that its denoising effect is better than that of some classical fixed kernel function-based CCTFDs, including the Margenau-Hill distribution, the Kirkwood-Rihaczek distribution, the Born-Jordan distribution, and the Page distribution. Under white Gaussian noise, our method outperforms the ordinary Wiener filter, while under colored noise, it performs comparably to the Wiener filter.
Appendix
In this section, we prove all the theoretical results in the letter.
V-A Proof of Eq. (10)

### v2

@@ 4 Numerical experiments
Red noise Margenau-Hill 2.38882.38882.3888 6.39986.39986.3998 3.63603.63603.6360 5.25625.25625.2562
Kirkwood-Rihaczek 1.77741.77741.7774 5.94565.94565.9456 3.48813.48813.4881 5.15595.15595.1559
Born-Jordan 4.01114.01114.0111 7.41677.41677.4167 4.38054.38054.3805 6.63486.63486.6348
Page 1.36061.36061.3606 5.98635.98635.9863 3.09253.09253.0925 3.86283.86283.8628
Adaptive radial Gaussian kernel 2.99442.99442.9944 6.62676.62676.6267 3.75523.75523.7552 3.40613.40613.4061
Wiener filter 11.883211.883211.8832 $\mathbf{16.7052}$ $\mathbf{13.1156}$ $\mathbf{11.9071}$
Adaptive CCTFD $\mathbf{20.4654}$ 16.149516.149516.1495 7.04967.04967.0496 7.20567.20567.2056
5 Conclusion
The convolution type of CCTFD time-frequency analysis method suitable for the adaptive filter denoising for additive noises jamming signals under the condition of low SNR was established, giving rise to the adaptive kernel function which minimizes the MSE in the WVD domain. The proposed adaptive CCTFD can automatically adjust its kernel function according to the change of signal to adapt to different signal characteristics. It turns out that its denoising effect is superior not only to some classical fixed kernel function-based CCTFDs, including the Margenau-Hill distribution, the Kirkwood-Rihaczek distribution, the Born-Jordan distribution, and the Page distribution, but also to the adaptive radial Gaussian kernel function-based CCTFD. Under white Gaussian noise, our method outperforms the ordinary Wiener filter, while under colored noise, it performs comparably to the Wiener filter.
[…]
Acknowledgements This work was supported by the National Natural Science Foundation of China (Grant No 61901223), the Jiangsu Planned Projects for Postdoctoral Research Funds (Grant No 2021K205B), the Foundation of Key Laboratory of System Control and Information Processing, Ministry of Education (Grant No Scip20240121), and the Foundation of Key Laboratory of Computational Science and Application of Hainan Province (Grant No JSKX202401).
Appendix A Proof of Eq. (10)